# Improving robustness to corruptions with multiplicative weight perturbations

### v2

@@ 5 Related works
Data augmentations aim at enhancing robustness include AugMix (Hendrycks et al., 2019), which combines common image transformations; Patch Gaussian (Lopes et al., 2019), which applies Gaussian noise to square patches; ANT (Rusak et al., 2020), which uses adversarially learned noise distributions for augmentation; and AutoAugment (Cubuk et al., 2018), which learns augmentation policies directly from the training data. These methods have been demonstrated to improve robustness to the corruptions in ImageNet-C (Hendrycks and Dietterich, 2019). Mintun et al. (2021) attribute the success of these methods to the fact that they generate augmented images perceptually similar to the corruptions in ImageNet-C and propose ImageNet-$\bar{\text{C}}$, a test set of 10 new corruptions that are challenging to models trained by these augmentation methods.
@@ Acknowledgments
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI and decision no. 359567, 345604 and 341763), ELISE Networks of Excellence Centres (EU Horizon: 2020 grant agreement 951847) and UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). We acknowledge the computational resources provided by Aalto Science-IT project and CSC–IT Center for Science, Finland.
@@ Appendix A Proof of Lemma 1
$||\nabla_{\mathbf{x}^{(h)}_{\mathbf{g}}}\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)-\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)||_{2}\leq L_{h}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}$ (23)
@@
$\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)\leq\ell(\boldsymbol{\omega},\mathbf{x},y)+\left\langle\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\right\rangle+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}$ (24)
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq\left\langle\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\right\rangle+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}$ (25)
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)},\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{L_{h}}{2}||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}$ (26)
@@
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}=||\mathbf{x}^{(0)}_{\mathbf{g}}-\mathbf{x}^{(0)}||_{2}^{2}=||\mathbf{g}(\mathbf{x})-\mathbf{x}||_{2}^{2}\leq M$ (27)
@@
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}=||\mathbf{x}^{(1)}_{\mathbf{g}}-\mathbf{x}^{(1)}||_{2}^{2}\leq||\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}$ (28)
Here we assume that the activate $\boldsymbol{\sigma}$ satisfies $||\boldsymbol{\sigma}(\mathbf{x})-\boldsymbol{\sigma}(\mathbf{y})||_{2}\leq||\mathbf{x}-\mathbf{y}||_{2}$, which is true for modern activation functions such as ReLU. Since $||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})$ such that:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}=||\mathbf{x}^{(1)}_{\mathbf{g}}-\mathbf{x}^{(1)}||_{2}^{2}\leq||\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})}{2}||\mathbf{W}^{(1)}||_{F}^{2}$ (29)
where $||\cdot||_{F}$ denotes the Frobenius norm. Similarly, as we have proven that $||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})$ such that:
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(2)}||_{2}^{2}=||\mathbf{x}^{(2)}_{\mathbf{g}}-\mathbf{x}^{(2)}||_{2}^{2}\leq||\mathbf{W}^{(2)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})}{2}||\mathbf{W}^{(2)}||_{F}^{2}$ (30)
@@
$||\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}||_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (31)
@@ Appendix B Proof of Theorem 1
$\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}_{k},\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{\hat{C}_{\mathbf{g}}^{(h)}}{2}||\mathbf{W}^{(h)}||_{F}^{2}$ (33)
@@
$\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h-1)}_{k},\mathbf{W}^{(h)}\right\rangle}_{F}\\ +\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (34)
@@
$\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}{\left\langle\frac{1}{N}\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h-1)}_{k},\mathbf{W}^{(h)}\right\rangle}_{F}+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (35)
$=\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\sum_{h=1}^{H}{\left\langle\nabla_{\mathbf{W}^{(h)}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S}),\mathbf{W}^{(h)}\circ\boldsymbol{\xi}^{(h)}(\mathbf{g})\right\rangle}_{F}+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (36)
$\leq\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi}(\mathbf{g});\mathcal{S})+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}=\mathcal{L}(\boldsymbol{\omega}\circ\left(1+\boldsymbol{\xi}(\mathbf{g})\right);\mathcal{S})+\frac{C_{\mathbf{g}}}{2}||\boldsymbol{\omega}||_{F}^{2}$ (37)
@@ Appendix E Corruption datasets
ImageNet-$\overline{\mathrm{C}}$ (Mintun et al., 2021)
This dataset contains the corrupted versions of the ImageNet validation set, as the labels of the true ImageNet test set was never released. It contains 10 types of corruption, each divided into 5 levels of severity. The types of corruption in ImageNet-$\overline{\mathrm{C}}$ differ from those in ImageNet-C.

### v3

@@ 5 Related works
Data augmentations aim at enhancing robustness include AugMix (Hendrycks et al., 2019), which combines common image transformations; Patch Gaussian (Lopes et al., 2019), which applies Gaussian noise to square patches; ANT (Rusak et al., 2020), which uses adversarially learned noise distributions for augmentation; and AutoAugment (Cubuk et al., 2018), which learns augmentation policies directly from the training data. These methods have been demonstrated to improve robustness to the corruptions in ImageNet-C (Hendrycks and Dietterich, 2019). Mintun et al. (2021) attribute the success of these methods to the fact that they generate augmented images perceptually similar to the corruptions in ImageNet-C and propose ImageNet-$\overline{\text{C}}$, a test set of 10 new corruptions that are challenging to models trained by these augmentation methods.
@@ Acknowledgments
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence FCAI and decision no. 359567, 345604 and 341763), ELISE Networks of Excellence Centres (EU Horizon: 2020 grant agreement 951847) and UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1). We acknowledge the computational resources provided by Aalto Science-IT project and CSC-IT Center for Science, Finland.
@@ Appendix A Proof of Lemma 1
$\|\nabla_{\mathbf{x}^{(h)}_{\mathbf{g}}}\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)-\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\|_{2}\leq L_{h}\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\|_{2}$ (23)
@@
$\ell(\boldsymbol{\omega},\mathbf{x}_{\mathbf{g}},y)\leq\ell(\boldsymbol{\omega},\mathbf{x},y)+\left\langle\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\right\rangle+\frac{L_{h}}{2}\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\|_{2}^{2}$ (24)
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq\left\langle\nabla_{\mathbf{x}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x},y),\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\right\rangle+\frac{L_{h}}{2}\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\|_{2}^{2}$ (25)
@@
$\boldsymbol{\delta}_{\mathbf{g}}\ell(\boldsymbol{\omega},\mathbf{x},y)\leq{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x},y)\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)},\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{L_{h}}{2}\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\|_{2}^{2}$ (26)
@@
$\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}\|_{2}^{2}=\|\mathbf{x}^{(0)}_{\mathbf{g}}-\mathbf{x}^{(0)}\|_{2}^{2}=\|\mathbf{g}(\mathbf{x})-\mathbf{x}\|_{2}^{2}\leq M$ (27)
@@
$\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}\|_{2}^{2}=\|\mathbf{x}^{(1)}_{\mathbf{g}}-\mathbf{x}^{(1)}\|_{2}^{2}\leq\|\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}\|_{2}^{2}$ (28)
Here we assume that the activate $\boldsymbol{\sigma}$ satisfies $\|\boldsymbol{\sigma}(\mathbf{x})-\boldsymbol{\sigma}(\mathbf{y})\|_{2}\leq\|\mathbf{x}-\mathbf{y}\|_{2}$, which is true for modern activation functions such as ReLU. Since $\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}\|_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})$ such that:
$\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}\|_{2}^{2}=\|\mathbf{x}^{(1)}_{\mathbf{g}}-\mathbf{x}^{(1)}\|_{2}^{2}\leq\|\mathbf{W}^{(1)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(0)}\|_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(1)}(\mathbf{x})}{2}\|\mathbf{W}^{(1)}\|_{F}^{2}$ (29)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. Similarly, as we have proven that $\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}\|_{2}^{2}$ is bounded, there exists a constant $\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})$ such that:
$\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(2)}\|_{2}^{2}=\|\mathbf{x}^{(2)}_{\mathbf{g}}-\mathbf{x}^{(2)}\|_{2}^{2}\leq\|\mathbf{W}^{(2)}\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(1)}\|_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(2)}(\mathbf{x})}{2}\|\mathbf{W}^{(2)}\|_{F}^{2}$ (30)
@@
$\|\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}\|_{2}^{2}\leq\frac{\hat{C}_{\mathbf{g}}^{(h)}(\mathbf{x})}{2}\|\mathbf{W}^{(h)}\|_{F}^{2}$ (31)
@@ Appendix B Proof of Theorem 1
$\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h+1)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h)}_{k},\mathbf{W}^{(h+1)}\right\rangle}_{F}+\frac{\hat{C}_{\mathbf{g}}^{(h)}}{2}\|\mathbf{W}^{(h)}\|_{F}^{2}$ (33)
@@
$\mathcal{L}(\boldsymbol{\omega};\mathbf{g}(\mathcal{S}))\leq\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}\frac{1}{N}\sum_{k=1}^{N}{\left\langle\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h-1)}_{k},\mathbf{W}^{(h)}\right\rangle}_{F}\\ +\frac{C_{\mathbf{g}}}{2}\|\boldsymbol{\omega}\|_{F}^{2}$ (34)
@@
$\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\frac{1}{H}\sum_{h=1}^{H}{\left\langle\frac{1}{N}\sum_{k=1}^{N}\nabla_{\mathbf{z}^{(h)}}\ell(\boldsymbol{\omega},\mathbf{x}_{k},y_{k})\otimes\boldsymbol{\delta}_{\mathbf{g}}\mathbf{x}^{(h-1)}_{k},\mathbf{W}^{(h)}\right\rangle}_{F}+\frac{C_{\mathbf{g}}}{2}\|\boldsymbol{\omega}\|_{F}^{2}$ (35)
$=\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\sum_{h=1}^{H}{\left\langle\nabla_{\mathbf{W}^{(h)}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S}),\mathbf{W}^{(h)}\circ\boldsymbol{\xi}^{(h)}(\mathbf{g})\right\rangle}_{F}+\frac{C_{\mathbf{g}}}{2}\|\boldsymbol{\omega}\|_{F}^{2}$ (36)
$\leq\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi}(\mathbf{g});\mathcal{S})+\frac{C_{\mathbf{g}}}{2}\|\boldsymbol{\omega}\|_{F}^{2}=\mathcal{L}(\boldsymbol{\omega}\circ\left(1+\boldsymbol{\xi}(\mathbf{g})\right);\mathcal{S})+\frac{C_{\mathbf{g}}}{2}\|\boldsymbol{\omega}\|_{F}^{2}$ (37)
@@ Appendix E Corruption datasets
ImageNet-$\overline{\text{C}}$ (Mintun et al., 2021)
This dataset contains the corrupted versions of the ImageNet validation set, as the labels of the true ImageNet test set was never released. It contains 10 types of corruption, each divided into 5 levels of severity. The types of corruption in ImageNet-$\overline{\text{C}}$ differ from those in ImageNet-C.
@@
ImageNet-Hard (Taesiri et al., 2023)
This dataset comprises an array of challenging images, curated from several validation datasets of ImageNet.